# On the Element-Wise Representation and Reasoning in Zero-Shot Image Recognition: A Systematic Survey

### v2

@@ VI Technical Details
[ZSP] means zero-shot prediction. [TTA] means test-time adaptation. [UA] means unsupervised adaptation. [VEncoder] means visual encoder. [TEncoder] means text encoder. [ViT-B]: ViT-B/32, ViT-B/16. [ViT-L]: ViT-L/14, ViT-L/14@336px. [Statistic] means the mean and variance of pretrained dataset. [Distribution] means the label distribution.
VI-A Dataset
We list in Table I the image recognition evaluation datasets commonly used for the three tasks. There are 8 datasets used for zero-shot object recognition, including fine-grained datasets such as CUB [9] which is a bird dataset, SUN [10] which is a scene dataset, and coarse-grained datasets such as AWA [154] which is a dataset of various types of animals. Most of the datasets contain detailed attribute annotations. Notice that Flowers102 [152] and NABirds [153] do not have attribute annotations. Flowers102 contains textual descriptions, i.e., each image is described by several sentences. NABirds has region annotations, which mark the parts of the bird, and are typically used in visual component analysis methods.
There are 6 commonly used datasets for zero-shot compositional recognition. The columen [Class/Attr.] indicates the number of objects and the number of states they contain. It can be seen that their compositional sizes cover different orders of magnitude, ranging from $3/8$ (AO-CLEVr [42]), which means up to 24 compositions, to $674/413$ (C-GQA [111]), which means 278,362 compositions.
We show 15 commonly used evaluation datasets for foundation model-based open-world recognition that cover different categories and different data sizes. ImageNet [1] includes data with up to 1000 categories, while EuroSAT [167] and CIFAR-10 [168] have only 10 categories. The data types are also diverse, including food [162], flowers [152], pets [165], transportation [163, 164], etc. They are also fine-grained datasets. In addition, DTD [166] is a texture dataset and EuroSAT [167] is a remote sensing dataset.
VI-B Implement Details
To facilitate understanding and referring readers to related work, we provide more details of the latest developments in Tables II and III. Table II shows the implementation details of the related studies of ZSOR and ZSCR, including the backbone, auxiliary, dataset, and code addresses. It can be seen that commonly used backbones include ResNet [175], VGG [176], GoogleNet [177], AlexNet [178], ViT [66], and CLIP [29]. [Auxiliary] refers to additional information such as region-level annotations. We provide the dataset and code address to facilitate access to the performance of the method as well as its implementation. Finally, we label the study with up to two keywords based on the specific technique used.
Table III shows more details of the related work in foundation model-based open-world recognition. We list their visual and text encoders, as well as auxiliary information. GPT-3 [133] and Diffusion [137] are the exotic models. [Statistic] denotes the mean and variance of the pre-training data. [Distribution] denotes the prior distribution of labels. We also provide their code addresses for quick reference. Since FBOR is at a young stage with less related work, it is not easy to categorize. To better understand the focus of each work, we summarize their contributions in a single sentence.
VI-C Shared Technologies
@@
Observing Tables II and III, we can identify some techniques that are common across the three tasks. We hope to discuss these techniques to reveal more commonalities between several tasks. Attention is the most commonly used technique, distributed across almost every sub-direction in the three tasks. This suggests that attention techniques are extremely generalizable and can be applied in different task goals and with different research motivations. In ZSOR, attention is used to localize key visual regions [36, 37], model visual-text correlations [39, 79, 40], and localize key attributes [89]. In ZSCR, attention is used to model object and state correlations [98, 101, 103], and enhance representations [109, 110]. In FBOR, attention is used to capture correlations between images and text [34]. Prototype learning is another more generalized technique. In ZSOR, prototypes are used to represent attributes [83, 85]. In ZSCR, prototypes are used to represent objects and states [94, 102]. In FBOR, prototypes are used to replace the embedding of prompts in deep space [148]. Graph and contrast learning are the more commonly used techniques in ZSOR and ZSCR. Graph is used to model the topological relationships between visual regions in ZSOR [70, 38], while in ZSCR it is used to capture the logical relationships between individual objects and states [111, 112]. Contrast learning is used in ZSOR to optimize attribute embeddings [77], while in ZSCR it is used to untangle and model semantic relations [49, 115]. In FBOR, thanks to the robust zero-shot performance of VLMs, fine-tuning prompts becomes a popular approach [132, 140]. Similarly, in ZSCR, some methods with CLIP as the backbone introduce prompt learning [102, 113], which effectively improves the performance of the model.
VII Application
Zero-shot learning is of great practical significance due to its radical freeing up of the pressure of image data collection. Meanwhile, element-wise zero-shot learning provides an intuitive humanoid study pattern which is suitable for a variety of reasoning tasks. In recent years, besides image recognition, element-wise zero-shot techniques have been applied in a wide range of areas, and they promote each other and develop synergistically.
@@ VII Application
Language and Text. Language, as one of the main media for human beings, has always received as much attention from researchers as images. Language research is also under pressure to label data. Therefore, zero-shot generalization ability is a long standing direction of attack [179, 180, 181]. In recent years, large language models have been developed rapidly [182], and they can even talk with humans fluently. This shows that large language models have powerful zero-shot reasoning ability. Meanwhile, some studies [183] propose fine-tuning schemes for language models to improve the zero-shot performance. In some specific tasks, zero-shot techniques have been widely applied to solve the problem of insufficient data samples, such as machine translation [184], question answering [185], etc. In addition, zero-shot techniques also play an important role in text analysis. Text analysis is the classification of text, such as sentiment classification [186], topic classification [187, 188], etc., which is extremely helpful for research in sociology, psychology and other disciplines. The content of text is quite complex, and the text elaborating the same topic may have multiple forms of expression. Obviously, the collection and labeling of samples is a huge obstacle, and zero-shot learning offers a promising solution for it. Moreover, zero-shot learning has been intensively studied in various fields such as named entity recognition [189] and information retrieval [190].
@@
Scene Understanding. Scene understanding is a topic based on image recognition but more challenging. Instead of assigning a label to an entire image, it usually identifies and localizes multiple elements in an image. For example, multi-label recognition task [191, 192] requires simultaneous recognition of multiple objects in an image. In contrast, object detection [193, 194, 195] and semantic segmentation [196, 197, 198] are two more complex and practical tasks. Object detection requires identifying and localizing multiple objects in an image, while semantic segmentation is a pixel-level classification task that requires assigning labels to each pixel according to semantic information. Scene understanding is an important research topic, which is not only the basis for AI to understand the world, but also a key part of deployment to industrial scenarios. Related technologies are needed in a variety of scenarios such as autonomous driving, medical diagnostic assistance, security monitoring, retail, and so on. However, they also require more detailed annotation and larger data for training than image recognition. Therefore, incorporating zero-shot learning into scene understanding has become an attractive direction in recent years. For example, zero-shot object detection [199, 200, 201] and zero-shot semantic segmentation [202, 203, 204] are dedicated to the deployment of models in the open world.
@@
3D Vision. With the increasing maturity of 3D sensors, the cost of acquiring 3D vision data has dropped significantly [205]. Compared with traditional 2D images, 3D visual data contains important depth information, and therefore can provide richer semantic information. 3D vision data has a wide range of applications, covering automatic driving, medical diagnosis, remote sensing, etc. Correspondingly, research on 3D vision data has also received more and more attention. However, compared with 2D data, the acquisition and labeling cost of 3D data is more expensive. Therefore, many researches integrate the concept of zero-shot into 3D tasks, such as point cloud classification [206, 207, 208], object detection [209], and semantic segmentation [210, 211]. In addition to semantic recognition tasks, some studies are also exploring the application of zero-shot learning to tasks such as generation [212], style transfer [213], and synthesis [214, 215]. There are also some studies that have explored in-depth on special tasks, such as human pose estimation [216], shape correspondence [217], and so on.
@@
Other Tasks. Zero-shot learning offers promising solutions for many specific application scenarios. For example, medical and remote sensing images are much more difficult to collect than natural images due to sample scarcity and data protection. Therefore, the research on zero-shot medical image [218, 219] recognition and remote sensing image [220, 221] recognition has been developed rapidly in recent years. Anomaly detection [222] plays a key role in securing many industrial fields. However, the number and type of anomaly samples are often insufficient. Some studies resort to zero-shot learning to solve this problem. In addition, zero-shot techniques have been applied in many special tasks to alleviate the low-data problem as well as to enhance the generalization ability of the model, such as action recognition [223], image retrieval [224, 225], and video classification [226].
VIII Future Directions
@@
VIII-1 Continual Zero-Shot Learning
Existing studies are based on closed systems to design solutions. However, in reality there are a large number of growing systems with dynamically changing environments and goals. For example, for zero-shot object recognition, attributes and categories may be progressively inflated. Early in the observation phase, fewer attributes are accessible, followed by a steady stream of new attribute annotations added. Similarly, the detected categories may increase over time forward. It is difficult for existing schemes to handle such dynamic recognition tasks, and continual learning has come a long way in this regard [227]. However, simply transplanting techniques from continual learning to zero-shot recognition tasks suffers from deviation. Therefore, fusing continual learning and zero-shot learning will further increase the value of applications to real-world scenarios.
VIII-2 Multi-label Compositional Recognition
Zero-shot compositional recognition allows the model to not only learn to recognize objects, but also to describe their state. However, current research is limited to single object and single state. A robust compositional recognition system should have the ability to recognize multiple objects as well as multiple states. For a single object, the recognition system should describe its state from multiple perspectives to provide more detailed information. For multiple objects, their individual states as well as their interactions, relationships, etc. should be described. Expanding from single label to multiple labels is a tough choice, but it is an inescapable step towards building a smarter recognition system.
VIII-3 Train VLM with Downstream Fine-Grained Annotations
The power of VLMs in the open world has been demonstrated by the fact that it requires only the names of the categories in the downstream task to accomplish zero-shot inference. However, the performance of VLMs in tasks with fine-grained annotations has not been explored. If the downstream tasks have more refined element-wise annotations, such as attribute texts or state descriptions, can VLMs fully utilize this information to obtain better zero-shot recognition performance. There are a number of issues that deserve deeper investigation, including domain shift between pre-trained text embeddings of VLMs and downstream text, fine-grained visual-text matching, catastrophic forgetting, and so on.
VIII-4 Annotation Cost and Quality
@@ VIII-5 Deployment Cost
Compared to class-wise semantic modeling, many element-wise schemes have to process a higher density of information, which introduces a more luxurious deployment cost. Such cost is reflected in bloated network structures and high computational complexity (Note that we discuss the deployment phase, excluding the training phase). As a result, they are unfriendly to edge tasks and mini-endpoints, which have to trade off performance and memory. However, they can be more favorable to a scenario associated with resource-constrained devices due to the low or even zero data requirements. Such a scenario can also well align with ubiquitous devices and data in real-world applications. Therefore, it is promising to investigate on-device-friendly element-wise algorithms.
VIII-6 Unified Framework
@@
Currently, zero-shot research for all directions is booming. They share many techniques such as attention, graph modeling, and so on. However, barriers between the them still exist due to the different goals and contexts of the tasks. A technique and framework that is versatile across all fields is lacking. But the significance of such a unified framework for both academia and application is immense. The quest for a unified framework means exploring more commonalities across zero-shot tasks, and thinking about the operational mechanisms of the entire realm from the ground-up principles. From a realistic point of view, a mature intelligence should be able to handle multiple tasks at the same time.
VIII-7 Robust Representation Learning
@@
Domain shift is a crucial hindrance to the generalization of zero-shot techniques to the open world, and one of its solutions is to learn robust feature representation. Although many methods have been proposed to mitigate the effects of domain shift, quantitative tests of the resistance to interference are lacking, and thus the robustness of existing representation methods in real-world environments remains unknowable. Recently Shafiee et al. [228] introduce adversarial attacks into zero-shot models and propose a framework for combining attributes to generate perturbations that effectively degrade the performance of zero-shot models. Attack and defense of models are two topics that have coexisted for a long time. Attackers seek to weaken the model at minimal cost, while defenders need to resist attacks. They learn from each other and progress continuously. Very often, the defender must learn more robust representation in order to resist perturbations. Thus, for zero-shot learning, it provides a new perspective to combat domain shift.
VIII-8 Poor Theoretical Foundation
@@
The development of zero-shot learning is established on the beautiful hypothesis that deep neural networks can reason logically like humans. As mentioned earlier, we can learn to identify basic elements and recognize new concepts by combining and reasoning about them. Nevertheless, there are not many solid theories on the compatibility between human reasoning and machine inductive ability so far, leading to a lack of explainability. Meanwhile, some flaws also challenge the plausibility of the hypothesis, such as the correspondence between abstract attributes and vision. Rigorous theoretical guidance is at the helm of a field moving forward, and it is of great prospective to dive into the mysterious black box in the future.
IX Conclusion

### v3

@@ 6.2 Implementation
APPLe [137] arXiv 2023 ViT-L/14 Transformer GPT-3 Propose prompt prototype learning by initializing from description sentences. -
@@
[ZSP] means zero-shot prediction. [TTA] means test-time adaptation. [UA] means unsupervised adaptation. [VEncoder] means visual encoder. [TEncoder] means text encoder. [ViT-B]: ViT-B/32, ViT-B/16. [ViT-L]: ViT-L/14, ViT-L/14@336px. [Statistic] means the mean and variance of pretrained dataset. [Distribution] means the label distribution. Click the [link] to check more information in the corresponding repository.
6.3 Shared Technologies
@@
It can be seen from Tables II, III, and IV that some techniques are common across the three tasks. We further discuss these shared techniques to reveal more commonalities between these tasks. In particular, attention is the most commonly used technique, distributed across almost every sub-direction in the three tasks. This suggests that attention techniques are extremely generalizable and can be applied to different task goals and with different research motivations. In ZSOR, attention is used to localize key visual regions [35, 36], model visual-text correlations [38, 81, 39], and localize key attributes [92]. In ZSCR, attention is used to model object and state correlations [101, 104, 107], and enhance representations [113, 114]. In FBOR, attention is used to capture correlations between images and text [33]. Prototype learning is another more generalized technique. In ZSOR, prototypes are used to represent attributes [85, 87]. In ZSCR, prototypes are used to represent objects and states [97, 105]. In FBOR, prototypes are used to replace the embedding of prompts in deep space [150]. Graph and contrast learning are the more commonly used techniques in ZSOR and ZSCR. The graph is used to model the topological relationships between visual regions in ZSOR [72, 37], while in ZSCR, it is used to capture the logical relationships between individual objects and states [115, 116]. Contrast learning is used in ZSOR to optimize attribute embeddings [79], while in ZSCR, it is used to untangle and model semantic relations [48, 119]. In FBOR, thanks to the robust zero-shot performance of VLMs, fine-tuning prompts become a popular approach [133, 142]. Similarly, in ZSCR, some methods with CLIP as the backbone introduce prompt learning [105, 117], which effectively improves the performance of the model.
7 Application
In recent years, besides image recognition, element-wise zero-shot techniques have been applied in a wide range of areas and developed synergistically. We ketch out some following related applications.
7.1 Language and Text
@@
Language, as one of the main media used by humans, has always received attention from researchers. Due to label pressures, zero-shot generalization is also a long-standing direction in language studies [180]. In recent years, large language models have been developed rapidly [181], and they can respond to human conversations fluently, showing their powerful zero-shot reasoning ability. Meanwhile, some studies [182] proposed fine-tuning schemes for language models to improve the zero-shot performance. In some specific tasks, zero-shot techniques have been widely applied to solve the problem of insufficient data samples, such as machine translation [183], question answering [184], etc. In addition, zero-shot techniques also play an important role in text analysis, such as sentiment classification [185], topic classification [186], etc., which is extremely helpful for research in sociology, psychology, and other disciplines. Meanwhile, the collection and labeling of text samples is a huge obstacle, and zero-shot generalization offers a promising solution [187]. Moreover, recent studies, such as named entity recognition [187] and information retrieval [188], also extensively rely on zero-shot ability.
7.2 Scene Understanding
@@
Scene understanding is a topic based on image recognition but is more challenging. Instead of assigning a label to an entire image, it usually identifies and localizes multiple elements in an image. For example, multi-label recognition task [189] requires simultaneous recognition of multiple objects in an image. In contrast, object detection [190, 191] and semantic segmentation [192, 193] are two more complex and practical tasks. Object detection requires identifying and localizing multiple objects in an image, while semantic segmentation is a pixel-level classification task that requires assigning labels to each pixel according to semantic information. Scene understanding is an important research topic, which is not only the basis for AI to understand the world but also a key part of deployment to industrial scenarios. Related technologies are needed in a variety of scenarios, such as autonomous driving, medical diagnostic assistance, security monitoring, retail, and so on. However, they also require more detailed annotation and larger data for training than image recognition. Therefore, incorporating zero-shot learning into scene understanding has become an attractive direction in recent years. For example, zero-shot object detection [194, 195, 196] and zero-shot semantic segmentation [197, 198, 199] are dedicated to the deployment of models in the open world.
7.3 3D Vision
@@
With the increasing maturity of 3D sensors, the cost of acquiring 3D vision data has dropped significantly [200]. Compared with traditional 2D images, 3D visual data contains important depth information and, therefore, can provide richer semantic information. 3D vision data has a wide range of applications, covering automatic driving, medical diagnosis, remote sensing, etc. Correspondingly, research on 3D vision data has also received more and more attention. However, compared with 2D data, the acquisition and labeling cost of 3D data is more expensive. Therefore, many researches integrate the concept of zero-shot into 3D tasks, such as point cloud classification [201, 202], object detection [203], and semantic segmentation [204]. In addition to semantic recognition tasks, some studies are also exploring the application of zero-shot learning to tasks such as generation [205], style transfer [206], and synthesis [207, 208]. There are also some studies that have explored in-depth on special tasks, such as human pose estimation [209], shape correspondence [210], and so on.
7.4 Ohters
@@
Zero-shot learning offers promising solutions for many specific application scenarios. For example, medical and remote sensing images are much more difficult to collect than natural images due to sample scarcity and data protection. Therefore, the research on zero-shot medical image [211] recognition and remote sensing image [212] recognition has been developed rapidly in recent years. Anomaly detection [213] plays a key role in securing many industrial fields. However, the number and type of anomaly samples are often insufficient. Some studies resort to zero-shot learning to solve this problem. In addition, zero-shot techniques have been applied in many special tasks to alleviate the low-data problem as well as to enhance the generalization ability of the model, such as action recognition [214], image retrieval [215], and video classification [216].
8 Future Directions
@@
8.1 Continual Zero-Shot Learning
Existing studies are based on closed systems to design solutions. However, in reality, there are a large number of growing systems with dynamically changing environments and goals. For example, for zero-shot object recognition, attributes and categories may be progressively inflated. Early in the observation phase, fewer attributes are accessible, followed by a steady stream of new attribute annotations added. Similarly, the detected categories may increase over time forward. It is difficult for existing schemes to handle such dynamic recognition tasks, and continual learning has come a long way in this regard [217]. However, simply transplanting techniques from continual learning to zero-shot recognition tasks suffers from deviation. Therefore, fusing continual learning and zero-shot learning will further increase the value of applications to real-world scenarios.
8.2 Multi-label Compositional Recognition
Zero-shot compositional recognition allows the model to not only learn to recognize objects but also to describe their state. However, current research is limited to a single object and single state. A robust compositional recognition system should have the ability to recognize multiple objects as well as multiple states. For a single object, the recognition system should describe its state from multiple perspectives to provide more detailed information [3]. For multiple objects, their individual states, as well as their interactions, relationships, etc., should be described [218]. Expanding from single labels to multiple labels is a tough choice, but it is an inescapable step toward building a smarter recognition system.
8.3 VLMs with Downstream Fine-Grained Annotations
The power of VLMs in the open world has been demonstrated by the fact that it requires only the names of the categories in the downstream task to accomplish zero-shot inference. Meanwhile, fine-grained annotations can improve the zero-shot inference performance of VLMs [32, 135]. What happens if VLMs are combined with downstream exclusive attribute annotations, such as attribute texts or state descriptions? There are a number of issues that deserve deeper investigation, including domain shift between pre-trained text embeddings of VLMs and downstream text, fine-grained visual-text matching, catastrophic forgetting, and so on.
8.4 Annotation Cost and Quality
@@ 8.5 Deployment Cost
Compared to class-wise semantic modeling, many element-wise schemes have to process a higher density of information, which introduces a more luxurious deployment cost. Such cost is reflected in bloated network structures and high computational complexity in the deployment phase. As a result, they are unfriendly to edge tasks and mini-endpoints, which have to trade off performance and memory [219]. However, they can be more favorable to a scenario associated with resource-constrained devices due to the low or even zero data requirements. Such a scenario can also well align with ubiquitous devices and data in real-world applications. Therefore, it is promising to investigate on-device-friendly element-wise algorithms.
8.6 Unified Framework
@@
Zero-shot studies share many common techniques, such as attention, graph modeling, and so on. However, barriers between them still exist due to the different goals and contexts of the tasks. A technique and framework that is versatile across all fields is lacking. However, the significance of such a unified framework for both academia and application is immense [220]. The quest for a unified framework means exploring more commonalities across zero-shot tasks and thinking about the operational mechanisms of the entire realm from the ground-up principles. From a realistic point of view, a mature intelligence should be able to handle multiple tasks at the same time.
8.7 Robust Representation Learning
@@
Despite some solutions have been proposed [91] to tackle with domain shift, there is a lack of quantitative assessments regarding their resistance to interference. Consequently, the robustness of existing representation methods in real-world environments remains uncertain. Recently, Shafiee et al. [221] introduce adversarial attack technique into zero-shot models. The interplay between model attacks and defenses has been a long-standing topic. Attackers aim to weaken models at minimal cost, while defenders strive to develop more robust representations to resist these perturbations. This dynamic process of learning from each other leads to continuous progress, offering a new perspective on combating domain shift.
8.8 Poor Theoretical Foundation
@@
The development of zero-shot learning is established on the beautiful hypothesis that deep neural networks can reason logically like humans [18, 22]. As mentioned, we can learn to identify basic elements and recognize new concepts by combining and reasoning about them. Nevertheless, there are not many solid theories on the compatibility between human reasoning and machine inductive ability, leading to a lack of explainability. Meanwhile, some flaws also challenge the plausibility of the hypothesis, such as the correspondence between abstract attributes and vision. Rigorous theoretical guidance is at the helm of a field moving forward, and it is of great prospective to dive into the mysterious black box in the future.
9 Conclusion